after striking the floor a certain ball rebounds 4/5 th of its height from which it has fallen.find total distance that it travel before coming to rest,if it is dropped from a height of 120 m
Answer by fractalier(6550)

This is an infinite sum...actually two sums...the down distances and the up distances...we find those sums by
S = a / (1-r)
For the down distances,
S(downs) = 120 / (1 - 4/5) = 600 m
For the bounces upward,
S(ups) = (4/5)(120) / (1 - 4/5) = 480 m
Total = 1080 m